A satellite going around Earth in an elliptic orbit has a speed of 10 km s−1 at the perigee which is at a distance of 227 km from the surface of the earth. Calculate the apogee distance and its speed at that point.


1
Expert's answer
2021-03-04T11:56:56-0500
"\frac{v^2_a}{2}-\frac{GM}{r_a}=\frac{v^2_p}{2}-\frac{GM}{r_p}\\
v_ar_a=v_pr_p\\
\frac{v^2_a}{2}-\frac{GM}{r_pv_p}v_a=\frac{v^2_p}{2}-\frac{GM}{r_p}\\"

"\frac{v^2_a}{2}-\frac{(6.67\cdot10^{-11})(5.97\cdot10^{24})}{227\cdot10^3(10^4)}v_a\\=\frac{(10^4)^2}{2}-\frac{(6.67\cdot10^{-11})(5.97\cdot10^{24})}{227\cdot10^3}\\"

"v_a=10\frac{km}{s}"

"r_a=227\ km"